OpenStudy (anonymous):

how do I solve x^2+8x all over x^2+10x+16 and put it in simplified form.

OpenStudy (agent0smith):

First step: factor out an x on the top\[\Large \frac{ x^2+8x }{ x^2+10x+16 } = \frac{x(x+8) }{ x^2+10x+16 }\] Now try to factor the bottom

OpenStudy (anonymous):

ok would the bottom be x(x+10)16

OpenStudy (anonymous):

no

OpenStudy (anonymous):

ok why did you do x+8 divide x+8)(x+2 instead of doing the gcf x+4)(x+2

OpenStudy (anonymous):

about x2+8x agent0smith did the job and replace it by x(x+8) , next x2+10x+16 is a quadratic equation , after calculating delta we got two solutions x1=-2 and x2=-8 ,we can factorize the quadratic equation like this (x-x1)(x-x2) since -2 and -8 are negative so it's (x+2)(x+8), then we will get x(x+8) /(x+2)(x+8) so we will subtract by x+8 ,we have x/x+2 the correct answer

OpenStudy (anonymous):

ok thank you guys